# & Télécom Paris, C2

### v5

@@ 4 Regularity $2$ and the small defect heuristic
Generating series reformulation of [10, Cor. 1.10].
Definition 5 ([26, Def. 1.5 & Th. 2.35])
@@
Definition 6 (cf. [10, §4A], after [23, Lect. 14])
@@
The square code distinguisher, and the filtration attack from [1] that extends it, work for codes ${\mathsf{C}}$ with ${\mathsf{C}}^{\langle 2\rangle}\varsubsetneq\mathbb{F}^{n}$, i.e. of regularity $>2$. This means that codes of regularity $2$ are hard to deal with under this approach. On the opposite, for us, codes of regularity $2$ are nice because Definition 6 means their Betti diagram is simple: it has only two nontrivial rows. Observe that most codes of interest have regularity $2$ (with the notable exception of self-dual codes).
@@
For $r=2$ this is proved directly. Now assume $r\geq 3$. Then the first two equalities are reformulations of Lemma 4. Now assume moreover ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$. Then $H_{{\mathsf{C}}}(z)=1+kz+n\frac{z^{2}}{1-z}$, so $B(z)=(1+kz)(1-z)^{k}+nz^{2}(1-z)^{k-1}$ by Proposition 6, thus $B_{r}=(-1)^{r-1}(r-1)\binom{k+1}{r}+(-1)^{r}n\binom{k-1}{r-2}=(-1)^{r-1}\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$. On the other hand, ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$ also means ${\mathsf{C}}$ has regularity $2$ in the sense of Definition 6, so (74) reduces to $B_{r}=(-1)^{r-1}\beta_{r-1,r}+(-1)^{r-2}\beta_{r-2,r}=(-1)^{r-1}\operatorname{ind}(\varphi_{r})$ and we conclude.
@@
The minimal resolution of ${\mathsf{C}}$ has length $k-1$, so $\beta_{i,j}=0$ for $i>k-1$, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
Regularity $2$ helps in computing the full Betti diagram of codes. See Propositions 10 and 11 in the Supplementary material for more.
@@
By (78) and Theorem 4.1, for any $[n,k]$-code ${\mathsf{C}}$ of regularity $2$ and for any $r\geq 2$ we have
$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}.$ (85)
We might interpret this lower bound as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$. Then, when ${\mathsf{C}}$ is random, it is quite natural to expect that no other syzygy will appear, or equivalently that (85) will be an equality, i.e. $\operatorname{def}(\varphi_{r})=0$.
@@
Another argument is the minimal resolution conjecture of [19]. It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq\frac{k(k+1)}{n}$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq\frac{k(k+1)}{n}$ or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$. However, two points require our attention:
@@
Concerning point 2, clearly, codes behave quite differently over a finite field and over an infinite field. For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not. This makes it desirable to investigate links between Betti numbers and distance properties of a code.
For any code ${\mathsf{C}}$, let us denote by $A_{i}({\mathsf{C}})$ the number of codewords of Hamming weight $i$ in ${\mathsf{C}}$.
@@
Let ${\mathsf{C}}$ be a $[n,k]_{q}$-code of regularity $2$, with minimum distance $d=\operatorname{d}_{\min}({\mathsf{C}})$ and dual minimum distance $d^{\perp}=\operatorname{d}_{\min}({\mathsf{C}}^{\perp})$.
@@
    Moreover if $d^{\perp}\leq\frac{k(k+1)}{n}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
@@
    Moreover if $k+1-d\geq\frac{k(k+1)}{n}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of ${\mathsf{C}}$ where this inequality is an equality).
It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq\frac{k(k+1)}{n}$ and for $\frac{k(k+1)}{n}\leq\ r\leq k+1-d$, when applicable (i.e. when these intervals are nonempty).
@@
Conversely, for a random code ${\mathsf{C}}$ among codes of given parameters $[n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min\left(d^{\perp},\frac{k(k+1)}{n}\right)$ and as $r\gg\max\left(\frac{k(k+1)}{n},k+1-d\right)$.
In particular $r_{\max}({\mathsf{C}})=\max\left(\left\lfloor\frac{k(k+1)}{n}\right\rfloor,k+1-d\right)$, or is very close to this value, with high probability.
Figures 13-14 in the Supplementary material illustrate various aspects of these Experimental facts.
Consequently, we postulate:
Heuristic 1
Fix a field cardinality $q$, assume $n$ is not too close to $k$ in order to stay away from the counterexamples to the minimal resolution conjecture, and $n<\binom{k+1}{2}$ in order to ensure regularity $2$. Then for random $[n,k]_{q}$-codes, with high probability:
  1. 1.
    if $d>k+1-\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=0$ for $r>\frac{k(k+1)}{n}$
  2. 2.
    if $d^{\perp}>\frac{k(k+1)}{n}$ we expect $\beta_{r-1,r}=\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$ for $r<\frac{k(k+1)}{n}$.
@@
5 The distinguisher
Basic version (without shortening)
We compared our new distinguisher (in this basic version) with the one from [7], on the very same examples taken from this reference: $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$ and $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$. Results can be found at the end of the Supplementary material. In both cases, we improve the distinguishability threshold. Interestingly, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{4,4,4}$, this improved threshold is the one given by (86) (with $r=r_{\max}^{*}=4$), i.e. the conditions in Heuristic 1 are always met. On the other hand, for $\operatorname{{\mathsf{Gop}}}^{\text{irr}}_{2,6,3}$, the conditions in Heuristic 1 are the limiting factor (but still improving over [7]).
Full version (with shortening)
We can also use shortened dual codes as input in our distinguisher. In general, shortening $s$ times a $[n,k]$-code ${\mathsf{C}}$ produces a $[n_{s},k_{s}]$-code ${\mathsf{C}}_{s}$ with $n_{s}=n-s$, $k_{s}=k-s$. Having smaller parameters will decrease the complexity. Moreover the (dual) rate $R_{s}=k_{s}/n_{s}$ also decreases when we shorten; as observed in Remark 2, this will help satisfying the conditions in Heuristic 1.
Example 1
Consider the class $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ of dual binary Goppa codes with $m=10$, $n=1024$, $t=10$, irreducible Goppa polynomial. Without shortening, these codes are out of reach of our distinguisher: we would have to run Algorithm 1 up to $r=10$ and do linear algebra in dimension roughly $10^{20}$, which is infeasible. On the other hand, we can effectively distinguish after shortening $40$ times: we experimentally find that $40$-shortened $\operatorname{{\mathsf{Gop}}}^{\text{irr},\perp}_{2,10,10}$ have $\beta_{3,4}=30$, while $\beta_{3,4}=0$ for random codes with the same parameters. Computational cost (several days each, on a workstation with several hundreds of GB of RAM) prevented the author to run more than a few tests, but no deviation was observed. In particular $\beta_{3,4}=0$ consistently in the random case confirms Heuristic 1 for these parameters.
Proposition 8
$r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ (87)
Set $k=mt$, and let ${\mathsf{C}}$ be either a random $[n,k]_{q}$-code, or an element of $\mathcal{T}_{q,m,n,t}$.
Choose an integer $s$ such that
$\frac{k_{s}(k_{s}+1)}{n_{s}}<r^{*}-s$ (88)
where $n_{s}=n-s$ and $k_{s}=k-s$.
  1. 1.
    Let $d_{s}\approx\operatorname{d}_{GV}(q,n_{s},k_{s})$ be the typical minimum distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$. Then, under part 1. of Heuristic 1, computing
    $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s}),$ (89)
    where ${\mathsf{C}}_{s}$ is any $s$-shortened subcode of ${\mathsf{C}}$, allows to distinguish whether ${\mathsf{C}}$ is a random code (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ w.h.p.) or an element of $\mathcal{T}_{q,m,n,t}$ (in which case $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$).
  2. 2.
    Moreover, let $d_{s}^{\perp}\approx\operatorname{d}_{GV}(q,n_{s},n_{s}-k_{s})$ be the typical dual distance of random $[n_{s},k_{s}]_{q}$-codes, and assume $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$. Then, under Heuristic 1, for large enough parameters the complexity of this distinguisher is dominated by the sum
    $\kappa=\sum_{i\leq\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1}\max\left(k_{s}\operatorname{ind}(\varphi_{i-1}),\binom{k_{s}+1}{2}\operatorname{ind}(\varphi_{i-2})\right)^{\omega}$ (90)
Proof
If ${\mathsf{C}}$ is random, so is ${\mathsf{C}}_{s}$. Then, as $d_{s}>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 1. of Heuristic 1 we indeed expect $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})=0$ for $r^{*}-s>\frac{k_{s}(k_{s}+1)}{n_{s}}$. On the other hand, for ${\mathsf{C}}$ in $\mathcal{T}_{q,m,n,t}$, the lower bound $r_{\max}({\mathsf{C}}_{s})\geq r^{*}-s$ precisely means $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})>0$.
Now using Algorithm 1 we obtain $\beta_{r^{*}-s-1,r^{*}-s}({\mathsf{C}}_{s})$ after computing the left kernel of the matrices $\mathbf{M}_{i}$ for $2\leq i\leq r^{*}-s$. The contributions for $i<4$ are negligible. For $i\geq 4$, the matrix $\mathbf{M}_{i}$ has size $k_{s}\beta_{i-2,i-1}({\mathsf{C}}_{s})\times\binom{k_{s}+1}{2}\beta_{i-3,i-2}({\mathsf{C}}_{s})$ and the contributions for $i>\left\lfloor\frac{k_{s}(k_{s}+1)}{n_{s}}\right\rfloor+1$ are negligible because of part 1. Then, as $d_{s}^{\perp}>\frac{k_{s}(k_{s}+1)}{n_{s}}$, under part 2. of Heuristic 1 we expect $\beta_{i-1,i}({\mathsf{C}}_{s})=\operatorname{ind}(\varphi_{i})$ for $2\leq i\leq\frac{k_{s}(k_{s}+1)}{n_{s}}$ for random ${\mathsf{C}}$; and we can assume this holds also for ${\mathsf{C}}\in\mathcal{T}_{q,m,n,t}$, otherwise computing this dimension readily provides a distinguisher with lower complexity. From this we conclude.
Example 2
The following table shows how the distinguisher from Proposition 8 applies to Classic McEliece parameters. We take $r^{*}=\widehat{f}$ as given by Corollary 4, and then choose the maximum possible shortening order $s$ satisfying (88).
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr(n,m,t)&(3488,12,64)&(4608,13,96)&(6688,13,128)&(6960,13,119)&(8192,13,128)\\ \hline\cr r^{*}&427&683&939&867&939\\ \hline\cr s&377&568&816&769&848\\ \hline\cr r^{*}-s&50&115&123&98&91\\ \hline\cr[n_{s},k_{s}]&[3111,391]&[4040,680]&[5872,848]&[6191,778]&[7344,816]\\ \hline\cr\frac{k_{s}(k_{s}+1)}{{n_{s}}_{\phantom{0\!\!}}}&49.27&114.62&122.61&97.89&90.78\\ \hline\cr\operatorname{d}_{GV}&921&1069&1650&1828&2256\\ \hline\cr\operatorname{d}_{GV}^{\perp}&55&62&122&108&110\\ \hline\cr\kappa&{2^{528}}^{\phantom{i\!}}&(2^{1080})&(2^{1224})&2^{1030}&2^{997}\\ \hline\cr\end{array}$
We observe that the condition $\operatorname{d}_{GV}(n_{s},k_{s})>k_{s}+1-\frac{k_{s}(k_{s}+1)}{n_{s}}$ is always satisfied with a large margin, so we could choose a smaller $s$ and still be able to distinguish. Concerning the complexity estimate $\kappa$, it relies on the condition $\operatorname{d}_{GV}(n_{s},n_{s}-k_{s})>\frac{k_{s}(k_{s}+1)}{n_{s}}$. We see this condition is satisfied for the parameter sets $(3488,12,64)$, $(6960,13,119)$, $(8192,13,128)$, but for $(4608,13,96)$ and $(6688,13,128)$ it is not (then, it would still be possible to give a complexity estimate, using Lemma 5 instead of Heuristic 1). These complexities improve those from [7], although they remain practically unreacheable and well beyond security levels.
Asymptotics
  • $m=\lceil\log_{q}(n)\rceil=\log_{q}(n)+O(1)$
  • $k\approx Rn\;$ such that:
  • $t=\frac{k}{m}\;$ is an integer.
Lemma 11
We have
Proof
Direct consequence of:
  • $e\;=\;\lfloor\log_{q}(t-1)\rfloor\;=\;\log_{q}(n)-\log_{q}\log_{q}(n)+O(1)$
  • $f\;=\;(e+1)t-\frac{q^{e+1}-1}{q-1}\;=\;et+O(t)$.
In the case of binary Goppa codes, the $\widehat{f}$ from Corollary 4 improves the $f$ from Theorem 3.1. However one could show that both have the same asymptotics. So, in order to distinguish these codes from random codes, we will not care about the extra structure: treating Goppa codes as alternant codes will suffice. (Still, this could be used to distinguish Goppa from “general” alternant codes.)
Now, under Heuristic 1, we have:
Theorem 5.1
$\kappa=q^{\left(\omega\frac{R^{2}}{1-R}+o(1)\right)\frac{(\log_{q}\log_{q}(n))^{3}}{(\log_{q}(n))^{2}}n}$ (92)
where $\omega\approx 2.372$ is the exponent of linear algebra.
Proof
We use Proposition 8 with $\mathcal{T}=\operatorname{{\mathsf{Alt}}}^{\perp}$, and $r^{*}=f$ as provided by Theorem 3.1. The analysis will be easier if rephrased in terms of $k_{r^{*}}=k-r^{*}$ and $r=r^{*}-s$. First, Lemma 11 gives
$k_{r^{*}}\;=\;k-f\;\sim\;R\frac{\log_{q}\log_{q}(n)}{\log_{q}(n)}n.$ (93)
Let us then turn to condition (88). Fix an $\varepsilon>0$, and set
Moreover the shortened code ${\mathsf{C}}_{s}$ has rate $\frac{k_{s}}{n_{s}}=\frac{k_{r^{*}}+r}{n-k+k_{r^{*}}+r}=o(1)$, so by Remark 2 both conditions in Heuristic 1 are satisfied. Thus Proposition 8 applies.

### v6

@@ 4 Regularity $2$ and the small defect heuristic
Generating series reformulation of [13, Cor. 1.10].
Definition 6([36, Def. 1.5 & Th. 2.35])
@@
Definition 7(cf. [13, §4A], after [31, Lect. 14])
@@
The square code distinguisher, and the filtration attack from [2] that extends it, work for codes $\mathsf{C}$} with ${\mathsf{C}}^{\langle 2\rangle}\varsubsetneq\mathbb{F}^{n}$, i.e. of regularity >2>2. This means that codes of regularity $2$ are hard to deal with under this approach. On the opposite, for us, codes of regularity $2$ are nice because Definition 7 means their Betti diagram is simple: it has only two nontrivial rows. Observe that most codes of interest have regularity $2$ (with the notable exception of self-dual codes).
@@
For $r=2$ this is proved directly. Now assume $r\geq 3$. Then the first two equalities are reformulations of Lemma 4. Now assume moreover ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$. Then $H_{{\mathsf{C}}}(z)=1+kz+n\frac{z^{2}}{1-z}$, so $B(z)=(1+kz)(1-z)^{k}+nz^{2}(1-z)^{k-1}$ by Proposition 8, thus $B_{r}=(-1)^{r-1}(r-1)\binom{k+1}{r}+(-1)^{r}n\binom{k-1}{r-2}=(-1)^{r-1}\left(\frac{k(k+1)}{r}-n\right)\binom{k-1}{r-2}$. On the other hand, ${\mathsf{C}}^{\langle 2\rangle}=\mathbb{F}^{n}$ also means $\mathsf{C}$} has regularity 22 in the sense of Definition 7, so (96) reduces to $B_{r}=(-1)^{r-1}\beta_{r-1,r}+(-1)^{r-2}\beta_{r-2,r}=(-1)^{r-1}\operatorname{ind}(\varphi_{r})$ and we conclude.
@@
The minimal resolution of $\mathsf{C}$} has length k1k-1, so $\beta_{i,j}=0$ for i>k1i>k-1, hence $\beta_{k-1,k+1}=-\operatorname{ind}(\varphi_{k+1})=n-k$.
Lemma 13([19, Th. 1], reformulated)
Proposition 10
$\beta_{r-1,r}=\frac{(r-1)(k-r)}{k}\binom{k+1}{r}$ (107)
for $2\leq r\leq k-1$. In particular they satisfy the symmetry $\beta_{i,j}=\beta_{k-1-i,k+1-j}$.
Proof
Proposition 11
Let $\mathsf{C}$} be a [2k1,k][2k-1,k] $\operatorname{{\mathsf{GRS}}}$ code. Then the nonzero Betti numbers of $\mathsf{C}$} are:
  • $\beta_{0,0}=1$
  • $\beta_{r-1,r}=(r-1)\binom{k-1}{r}\;$ for $2\leq r\leq k-1$
Proof
Now set n=2k1n=2k-1.
Figures 10 and 10 illustrate Propositions 10 and $1$.
Figure 9: a $[9,8]$ parity or $\operatorname{{\mathsf{GRS}}}$ code
Figure 10: a $[15,8]$ $\operatorname{{\mathsf{GRS}}}$ code
Exercise 1
Compute the Betti diagram of $[n,k]$ $\operatorname{{\mathsf{GRS}}}$ codes for k+1<n<2k1k+1<n<2k-1.
@@
Because of (103), a necessary condition for a $[n,k]$-code $\mathsf{C}$} of regularity $2$ to have $\beta_{r-1,r}=0$ is that $\operatorname{ind}(\varphi_{r})\leq 0$. This immediately yields
$r_{\max}({\mathsf{C}})\geq\lceil r_{\operatorname{crit}}({\mathsf{C}})\rceil-1$ (108)
where we define the critical degree of $\mathsf{C}$}
$r_{\operatorname{crit}}({\mathsf{C}})=\frac{k(k+1)}{n}$ (109)
as the value of $r$ at which the formula (105) for $\operatorname{ind}(\varphi_{r})$ changes sign.
More precisely, (100) and Theorem 4.1 give the lower bound
$\beta_{r-1,r}({\mathsf{C}})\geq\left(\frac{k(k+1)}{r}-n\right)^{+}\binom{k-1}{r-2}$ (110)
that can be interpreted as the number of “contingent” syzygies that are forced to exist, solely because of the dimension difference between the source and target spaces of the map $\varphi_{r}$. Then, when $\mathsf{C}$} is random, it is quite natural to expect that no other syzygy will appear, or equivalently that (110) will be an equality, i.e. $\operatorname{def}(\varphi_{r})=0$.
@@
Another argument is the minimal resolution conjecture of [25]. It postulates that, over an infinite field, a Zariski-generic code satisfies $\beta_{r-1,r}({\mathsf{C}})=0$ for $r\geq r_{\operatorname{crit}}({\mathsf{C}})$ and $\beta_{r-2,r}({\mathsf{C}})=0$ for $r\leq r_{\operatorname{crit}}({\mathsf{C}})$, or equivalently, $\operatorname{def}(\varphi_{r})=0$ for all $r\geq 2$. However, two points require our attention:
@@
Concerning point 2, clearly, codes behave quite differently over a finite field and over an infinite field. For instance, generic codes over an infinite field are MDS for any parameter set, while over a finite field they clearly are not. This makes it desirable to investigate links between Betti numbers and distance properties of a code. At this moment the author mostly has only experimental results in this direction: Actually, experiments suggest much stronger results hold. For any code $\mathsf{C}$}, let us denote by $A_{i}({\mathsf{C}})$ the number of codewords of Hamming weight $i$ in $\mathsf{C}$}. Also recall $r_{\operatorname{crit}}=\frac{k(k+1)}{n}$.
@@
Let $\mathsf{C}$} be a $n,k]_{q}-$code of regularity $2$, with minimum distance $d=\operatorname{d}_{\min}({\mathsf{C}})$ and dual minimum distance $d^{\perp}=\operatorname{d}_{\min}({\mathsf{C}}^{\perp})$.
@@
    Moreover if $d^{\perp}\leq r_{\operatorname{crit}}$ we have $\beta_{d^{\perp}-2,d^{\perp}}({\mathsf{C}})\geq A_{d^{\perp}}({\mathsf{C}}^{\perp})/(q-1)$ (and there are numerous examples of $\mathsf{C}$} where this inequality is an equality).
@@
    Moreover if $k+1-d\geq r_{\operatorname{crit}}$ we have $\beta_{k-d,k+1-d}({\mathsf{C}})\geq A_{d}({\mathsf{C}})/(q-1)$ (and there are numerous examples of $\mathsf{C}$} where this inequality is an equality).
It follows that $\operatorname{def}(\varphi_{r})>0$ for $d^{\perp}\leq r\leq r_{\operatorname{crit}}$ and for $r_{\operatorname{crit}}\leq\ r\leq k+1-d$, when applicable (i.e. when these intervals are nonempty).
@@
Conversely, for a random code $\mathsf{C}$} among codes of given parameters $n,k,d,d^{\perp}]_{q}$, the probability that $\operatorname{def}(\varphi_{r})=0$ tends quickly to $1$ as $r\ll\min(d^{\perp},r_{\operatorname{crit}})$ and as $r\gg\max(r_{\operatorname{crit}},k+1-d)$.
In particular $r_{\max}({\mathsf{C}})=\max(\lfloor r_{\operatorname{crit}}\rfloor,k+1-d)$, or is very close to this value, with high probability.
Actually, as a modest first step, the author is able to prove the first assertion in point 2. of Experimental fact 1:
Proposition 12
Any $[n,k,d]$ code $\mathsf{C}$} has $r_{\max}({\mathsf{C}})\geq k+1-d$, i.e. $\beta_{r-1,r}({\mathsf{C}})>0$ for all $r\leq k+1-d$.
Proof
Let $\mathbf{c}_{1}$ be a codeword of minimum weight $\operatorname{w}(\mathbf{c}_{1})=d$. Let $S$ be the support of $\mathbf{c}_{1}$, and let ${\mathsf{C}}_{S}\subseteq{\mathsf{C}}$ be the subcode made of all codewords that vanish over $S$. Then $\dim({\mathsf{C}}_{S})\geq k-\#{S}=k-d$, so we can find kdk-d linearly independent codewords $\mathbf{c}_{2},\dots,\mathbf{c}_{k+1-d}$ in ${\mathsf{C}}_{S}$. Let then complete $\mathbf{c}_{1},\dots,\mathbf{c}_{k-d+1}$ into a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of $\mathsf{C}$}, and use this basis to define the graded evaluation map $S=\mathbb{F}[X_{1},\dots,X_{k}]\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$. By construction we have $\mathbf{c}_{1}\mathbf{c}_{2}=\cdots=\mathbf{c}_{1}\mathbf{c}_{k-d+1}=\mathbf{0}$, which means $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle\subseteq I_{2}({\mathsf{C}})$. By Corollary 2 we deduce $\beta_{r-1,r}({\mathsf{C}})\geq\beta_{r-1,r}(S/\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle)$ for all $r$, hence in particular $r_{\max}({\mathsf{C}})\geq r_{\max}(S/\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle)$.
However we have $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle=X_{1}\langle X_{2},\dots,X_{k+1-d}\rangle$, so the minimal resolution of $\langle X_{1}X_{2},\dots,X_{1}X_{k+1-d}\rangle$ can be identified with the Koszul complex on $X_{2},\dots,X_{k+1-d}$ (with degree shifted by $1$), so its $r_{\max}$ is k+1dk+1-d.
@@
The quantity $r_{\max}=\max\{r:\;\beta_{r-1,r}>0\}$, and the dual quantity $r_{\min}=\min\{r:\;\beta_{r-2,r}>0\}$ that implicitely appears together with it in these Experimental facts, are at the core of deep conjectures of Green and Green-Lazarsfeld, discussed e.g. in the last two chapters of [13] (with the notation therein, $r_{\max}=b(X)$ and $r_{\min}=a(X)+3$).
Let us now illustrate various aspects of these Experimental facts.
Figures 13-13 compare Betti diagrams for three codes of length $n=23$ and dimension $k=12$ (so $r_{\operatorname{crit}}\approx 6.78$) of quite different nature.
First, Figure 13 presents the Betti diagram as predicted by the minimal resolution conjecture, i.e. with all defects equal to zero. It turns out that the minimal resolution conjecture is true for these parameters, so generic $[23,12]$-codes over an infinite field actually have this Betti diagram. Such codes are MDS, i.e. they have $d=12$ and $d^{\perp}=13$, so the intervals $d^{\perp}\leq r\leq r_{\operatorname{crit}}$ and $r_{\operatorname{crit}}\leq\ r\leq k+1-d$ are empty, and Experimental fact 1 is trivially verified. Actually, codes with this Betti diagram can be found even over not-so-large finite fields, e.g. $q=5$ suffices (the author did not search for optimality). These codes are no longer MDS, but they still have $d,d^{\perp}\geq 7$ and again Experimental fact 1 is trivially verified. Moreover, the converse Experimental fact 2 is also verified as all defects are zero.